Growth factors for the population of Chattanooga in the past two years have been 8 and 12. Find the value of the geometric mean


1
Expert's answer
2021-02-24T07:03:20-0500

"\text{Population growth factors are indicated in numerical form (not in percentage).}"

"\text{Therefore, the geometric mean will be}"

"gm=\sqrt{x_1x_2}=\sqrt{8*12}=\sqrt{96}=4\sqrt{16}"

Answer: "\sqrt{96}"